A temperature of 41 degrees F is equivalent to 278.2 degrees K, while a temperature of 86 degrees F is equivalent to 303.2 degrees K. Find the equation converting Kelvin temperature to Fahrenheit. What is the Fahrenheit temperature when it is 180 degrees K?
Answer by Nate(3500)

C = 5(F - 32)/9
and
K = C + 273.15
so
K = 5(F - 32)/9 + 273.15
K - 273.15 = 5(F - 32)/9
(K - 273.15)(9/5) + 32 = F
(180 - 273.15)(9/5) + 32 = F = -135.67 That's really cold.